a man's age is 3 times the sum of the ages of his two sons.after 5years his age will be twice the sum of the ages of his two sons.find the age of the man.
Answer by mananth(16949)

let age of sons be x & y & father's age =z
z = 3(x+y)
z/3 = (x+y)
..
after 5 years
their ages will be
x+5, y+5, z+5
z+5=2(x+5+y+5)
z+5=2x+2y+20
z+5=2(x+y)+20
z+5-20=2(x+y)+20-20
z-15=2(x+y)
But (x+y)= z/3
z-15=2*z/3
multiply 3
3*(z-15)=2z
3z-45=2z
3z-2z=45
z=45 father's age